# Single-cone Dirac edge states on a lattice

### v1

@@ I Introduction
The outline of the paper is as follows. In the next section we show how a zero-current boundary condition can be incorporated into the generalized eigenvalue problem on a 2D square lattice. In Sec. III we use this to study the edge states that appear for the zigzag boundary condition, and for more general boundary conditions that interpolate between zigzag and infinite mass. On the surface of a topological insulator, the entire class of boundary conditions is realized at the interface with a magnetic insulator, dependent on the orientation of the magnetization [21]. We compare with two alternative discretization schemes in Sec. IV and conclude in Sec. V with a discussion of the limitations of our approach.
@@ II.2 Generalized eigenvalue problem on the lattice
Both operators ${\cal H}$ and ${\cal P}$ in Eq. (11) are Hermitian and ${\cal P}$ is also positive definite. These two properties are both needed to ensure that ${\cal H}\Psi=E{\cal P}\Psi$ has real eigenvalues $E$. Moreover, ${\cal H}$ and ${\cal P}$ are sparse matrices, only nearby sites on the lattice are coupled, which allows for an efficient calculation of the energy spectrum.
@@ IV.2 Staggered fermions
V Conclusion
@@ Acknowledgements.
Our tangent fermion project has benefited much from the insights of M. J. Pacholski. Discussions with A. R. Akhmerov are gratefully acknowledged. Financial support was received from the European Research Council (ERC), under the European Union’s Horizon 2020 research and innovation programme.

### v2

@@ I Introduction
The outline of the paper is as follows. In the next section we show how a zero-current boundary condition can be incorporated into the generalized eigenvalue problem on a 2D square lattice. In Sec. III we use this to study the edge states that appear for the zigzag boundary condition, and for more general boundary conditions that interpolate between zigzag and infinite mass. On the surface of a topological insulator, the entire class of boundary conditions is realized at the interface with a magnetic insulator, dependent on the orientation of the magnetization [21]. We compare with two alternative discretization schemes and with one lattice-free calculation in Secs. IV and V. We then conclude in Sec. VI with a discussion of the limitations of our approach.
@@ II.2 Generalized eigenvalue problem on the lattice
Both operators ${\cal H}$ and ${\cal P}$ in Eq. (11) are Hermitian and ${\cal P}$ is also positive definite, These two properties are both needed to ensure that ${\cal H}\Psi=E{\cal P}\Psi$ has real eigenvalues $E$. Moreover, ${\cal H}$ and ${\cal P}$ are sparse matrices, only nearby sites on the lattice are coupled, which allows for an efficient calculation of the energy spectrum.
@@ IV.2 Staggered fermions
V Comparison with a lattice-free calculation
A lattice-free numerical solution of the Dirac equation in a confined geometry is developed in Ref. 10. That work uses infinite-mass boundary conditions, with a nonzero mass $m$ of the Dirac fermions inside the system. Although in the main text we took $m=0$, we can readily include a mass in the tangent fermion discretization [16]: This enters as a term $m\sigma_{z}{\cal P}\Psi$ on the left-hand-side of Eq. (11).
Figure 5: First five positive eigenvalues of the dimensionless Dirac equation, $-i\sigma_{x}\partial\psi/\partial x-i\sigma_{y}\partial\psi/\partial y+m\sigma_{z}\psi=E\psi$, in a rectangle of unit area with infinite-mass boundary conditions. The mass $m=1$ inside the rectangle, the aspect ratio of the rectangle is varied. The plot compares our lattice calculation (data points, lattice constant $1/80$) with the lattice-free calculation of Ref. 10 (solid lines).
Figure 6: Same as Fig. 5, but now showing only the first eigenvalue as a function of the mass $m$, at fixed aspect ratio of the rectangle. For this plot the lattice constant is set at $1/68$.
We compare results for a rectangle in Figs. 5 and 6. The agreement is quite satisfactory.
VI Conclusion
@@ Acknowledgements.
Our tangent fermion project has benefited much from the insights of M. J. Pacholski. Discussions with A. R. Akhmerov are gratefully acknowledged. Financial support was received from the European Research Council (ERC), under the European Union’s Horizon 2020 research and innovation programme. The computer codes used for Figs. 5 and 6 are available at https://dx.doi.org/10.5281/zenodo.14883054